one pan pizza and two cheesburgers provide 2860 claories. two pan pizzas and one cheeseburger provide 2990 calories. find the caloric content of each item
Answer by mananth(16949)

one pan pizza and two cheesburgers provide 2860 claories. two pan pizzas and one cheeseburger provide 2990 calories. find the caloric content of each item
...
pan pizza --- x calories
cheeseburgers - y calories
..
x+2y=2860.....................1
2x+y=2990......................2
...
multiply equation 1 by -2
-2x-4y=-5720
add it to equation2
..
2x-2x+y-4y=2990-5720
-3y=-2730
/-3
y=-2730/-3
y=910 CALORIES PER CHEESEBURGER
plug value of y in1
x+1820=2860
-1820
x=1040 calories in pan pizza

m.annath@hotmail.ca